# Nonlinear Schrödinger-Poisson systems in dimension two: the zero mass case

### v1

@@ 4.1. Mountain pass level estimate in the critical case
$\begin{split}\int_{\mathbb{R}^{2}}\ln(b+|x|)\widetilde{u}_{n}^{p}\,{\rm e}^{C_{\varepsilon}\alpha^{\prime}\frac{pq^{\prime}}{p-2}\tilde{u}^{2}}\,{\rm d}x&\leq\int_{\{\widetilde{u}\leq\widetilde{u}_{n}\}}\ln(b+|x|)\widetilde{u}_{n}^{p}\,{\rm e}^{C_{\varepsilon}\alpha^{\prime}\frac{pq^{\prime}}{p-2}\tilde{u_{n}}^{2}}\,{\rm d}x\\ &\quad+\int_{\{\widetilde{u}>\widetilde{u}_{n}\}}\ln(b+|x|)\widetilde{u}^{p}\,{\rm e}^{C_{\varepsilon}\alpha^{\prime}\frac{pq^{\prime}}{p-2}\tilde{u}^{2}}\,{\rm d}x\leq C\end{split}$
by (5.28)-(5.29). One may argue similarly for the third term in (5.27). As a result, we obtained (5.26) which, together with (5.24)-(5.25), (5.22), (5.18), show that (5.17) holds, yielding (5.16). Hence, by (5.14), one finally infers that $\widetilde{u}_{n}\to\widetilde{u}$ in $X$. Since $\mathcal{I}$ is a $C^{1}$-functional, then $\widetilde{u}$ is a weak solution of (1.1), which is nontrivial thanks to (5.8). Since $\widetilde{u}_{n}\geq 0$, by Remark 1.6 the solution $\widetilde{u}$ is positive in $\mathbb{R}^{2}$.
@@
$\begin{split}\int_{\mathbb{R}^{2}}\frac{|\Phi_{u}(x)|}{1+|x|^{2+2s}}\,{\rm d}x&\leq\int_{\mathbb{R}^{2}}F(u(y))\left(\int_{\mathbb{R}^{2}}\left|\ln\frac{1}{|x-y|}\right|\frac{1}{1+|x|^{2+2s}}\,{\rm d}x\right)\,{\rm d}y\\ &\leq\int_{\mathbb{R}^{2}}F(u(y))\left(\int_{\{|x-y|>1\}}\frac{\ln|x-y|}{1+|x|^{2+2s}}\,{\rm d}x+\int_{\{|x-y|\leq 1\}}\ln\frac{1}{|x-y|}\,{\rm d}x\right)\,{\rm d}y\\ &\leq\|F(u(y))\|_{1}\int_{\mathbb{R}^{2}}\frac{\ln(1+|x|)}{1+|x|^{2+2s}}\,{\rm d}x+\|F(u(y))\|_{\ast,1}\int_{\mathbb{R}^{2}}\frac{\,{\rm d}x}{1+|x|^{2+2s}}\\ &\quad+\|\ln(\cdot)\|_{L^{1}(B_{1}(0))}\int_{\mathbb{R}^{2}}F(u(y))\,{\rm d}y<+\infty\end{split}$
@@
$\begin{split}\widetilde{w}_{u}(x)-\Phi_{u}(x)&=\int_{\mathbb{R}^{2}}\left(\ln\left(\frac{1+|y|}{|x-y|}\right)-\ln\left(\frac{1}{|x-y|}\right)\right)F(u(y))\,{\rm d}y\\ &=\int_{\mathbb{R}^{2}}\ln(1+|y|)F(u(y))\,{\rm d}y<\|F(u)\|_{\ast,1}<+\infty\,,\end{split}$
@@ Acknowledgements:
The authors are members of Gruppo Nazionale per l’Analisi Matematica, la Probabilità e le loro Applicazioni (GNAMPA) of the Instituto Nazionale di Alta Matematica (INdAM). Federico Bernini is partially supported by INdAM-GNAMPA Project 2024 titled Problemi spettrali e di ottimizzazione di forma: aspetti qualitative e stime quantitative. Giulio Romani and Cristina Tarsi are partially supported by INdAM-GNAMPA Project 2023 titled Interplay between parabolic and elliptic PDEs and INdAM-GNAMPA Project 2024 titled New perspectives on Choquard equation through PDEs with local sources.

### v2

@@ 4.1. Mountain pass level estimate in the critical case
×$\quad\times\left(C\int_{\{\widetilde{u}\leq 1\}}|\widetilde{u}_{n}-\widetilde{u}|^{r}+\int_{\{\widetilde{u}>1\}}|\widetilde{u}_{n}-\widetilde{u}|^{r}|\widetilde{u}|^{\kappa}{\rm e}^{C_{\varepsilon}\alpha s^{\prime}\,\widetilde{u}^{2}}\right)^{\frac{1}{s^{\prime}}}$
$\leq\text{o}_{n}(1)+\left(\text{o}_{n}(1)+\|\widetilde{u}_{n}-\widetilde{u}\|_{rq^{\prime}}^{r/s}\left(\int_{\mathbb{R}^{2}}|\widetilde{u}_{n}|^{\kappa q}{\rm e}^{(1+\varepsilon)\alpha sq\,\widetilde{u}_{n}^{2}}\right)^{\frac{1}{sq}}\right)$
×$\quad\times\left(\text{o}_{n}(1)+\|\widetilde{u}_{n}-\widetilde{u}\|_{rq^{\prime}}^{r/s^{\prime}}\left(\int_{\mathbb{R}^{2}}|\widetilde{u}|^{\kappa q}{\rm e}^{C_{\varepsilon}\alpha s^{\prime}\,{\widetilde{u}^{2}}}\right)^{\frac{1}{s^{\prime}q}}\right).$
Defining $v_{n}:=G(\widetilde{u}_{n})$, by Lemma 4.6 we can estimate the first remaining term as follows:
$\int_{\mathbb{R}^{2}}|\widetilde{u}_{n}|^{\kappa q}{\rm e}^{(1+\varepsilon)\alpha sq\,\widetilde{u}_{n}^{2}}$ $\leq\frac{1}{\tau^{\kappa/2}}\int_{\mathbb{R}^{2}}|v_{n}|^{\kappa q}{\rm e}^{(1+\varepsilon)\alpha sq\left(t_{\varepsilon}+\frac{v_{n}}{1-\varepsilon}\right)^{2}}$
$\leq\frac{1}{\tau^{\kappa/2}}{\rm e}^{(1+\varepsilon)\alpha sC_{\varepsilon}t_{\varepsilon}^{2}}\int_{\mathbb{R}^{2}}|v_{n}|^{\kappa q}{\rm e}^{\left(\frac{1+\varepsilon}{1-\varepsilon}\right)^{2}\alpha sqv_{n}^{2}}.$
Fixing $\varepsilon>0$ close to 00, $\alpha>\alpha_{0}$ close to $\alpha_{0}$, $s>1$ close to $1$, and $\kappa=p/q$, recalling (4.33), we can now use Theorem 2.2 and find that
$\int_{\mathbb{R}^{2}}|\widetilde{u}_{n}|^{\kappa q}{\rm e}^{(1+\varepsilon)\alpha sq\,\widetilde{u}_{n}^{2}}$ $\leq C\|v_{n}\|_{p}^{p}\leq C\|u_{n}\|_{p}^{p}\leq C$
since $\{\widetilde{u}_{n}\}_{n}$ is bounded in $L^{p}(\mathbb{R}^{2})$. On the other hand, the second integral on the right in (5) is independent on $n$, and therefore, for the above choices of the parameters it remains bounded again by Theorem 2.2. Consequently, combining (5) with (5) or (5.20), we get
(5.22) $\int_{\mathbb{R}^{2}}\big{|}\widetilde{f}(\widetilde{u}_{n})-\widetilde{f}(\widetilde{u}_{n}-\widetilde{u})\big{|}|\widetilde{u}_{n}-\widetilde{u}|=\text{o}_{n}(1)\,.$
It remains to prove that also the last term in (5.18) vanishes. To this aim, since $f\in C^{1}(\mathbb{R})$ and so is $\widetilde{f}$ (cf. (5.15)), by Lagrange’s theorem,
$\widetilde{f}(\widetilde{u}_{n})-\widetilde{f}(\widetilde{u}_{n}-\widetilde{u})=\widetilde{f}^{\prime}(w_{n})\widetilde{u}=f^{\prime}(|w_{n}|)\widetilde{u}\,,$
where $w_{n}:=\theta_{n}\widetilde{u}_{n}+(1-\theta_{n})\widetilde{u}$ with $\theta_{n}:\mathbb{R}^{2}\to[0,1]$. Therefore, splitting $\mathbb{R}^{2}=B_{R}\cup B_{R}^{c}$ for a fixed $R>0$, we obtain
$\int_{\mathbb{R}^{2}}\ln(b+|x|)$ $\left|\widetilde{f}(\widetilde{u}_{n})-\widetilde{f}(\widetilde{u}_{n}-\widetilde{u})\right||\widetilde{u}_{n}-\widetilde{u}|\,{\rm d}x$
(5.23) $\ \ \leq\ln(b+R)\!\int_{B_{R}}\!\big{|}\widetilde{f}(\widetilde{u}_{n})-\widetilde{f}(\widetilde{u}_{n}-\widetilde{u})\big{|}|\widetilde{u}_{n}-\widetilde{u}|\,{\rm d}x$
$\quad+\!\int_{B_{R}^{c}}\!\!\ln(b+|x|)|f^{\prime}(|w_{n}|)|\,\widetilde{u}\,|\widetilde{u}_{n}-\widetilde{u}|\,{\rm d}x$
$\leq\text{o}_{n}(1)+\int_{B_{R}^{c}}\!\!\ln(b+|x|)|f^{\prime}(|w_{n}|)|\,\widetilde{u}\,|\widetilde{u}_{n}-\widetilde{u}|\,{\rm d}x$
by (5.22). To estimate this last term, we make use of assumption ($f_{5}$) on $f^{\prime}$, which implies that
$f^{\prime}(t)\leq C_{1}|t|^{p-2}+C_{2}|t|^{r}{\rm e}^{\alpha t^{2}}$
for $r>0$ and $\alpha>\alpha_{0}$, and some constants $C_{1},\,C_{2}>0$. Therefore we may estimate as follows
(5.24) $\int_{B_{R}^{c}}\!\!\ln(b$ $+|x|)|f^{\prime}(|w_{n}|)|\,\widetilde{u}\,|\widetilde{u}_{n}-\widetilde{u}|\,{\rm d}x$
$\lesssim\int_{B_{R}^{c}}\!\!\ln(b+|x|)|w_{n}|^{p-2}\,\widetilde{u}|\widetilde{u}_{n}-\widetilde{u}|\,{\rm d}x+\int_{B_{R}^{c}}\!\!\ln(b+|x|)|w_{n}|^{r}{\rm e}^{\alpha w_{n}^{2}}\,\widetilde{u}|\widetilde{u}_{n}-\widetilde{u}|=:T_{1}+T_{2}\,,$
and consider separately the two integrals. By Hölder’s inequality, using $|w_{n}|\leq|\widetilde{u}_{n}|+|\widetilde{u}|$, and $\|\widetilde{u}_{n}\|_{\ast,p}\leq C$ independently of $n$, we first have
(5.25) $T_{1}$ $\lesssim\!\left(\int_{B_{R}^{c}}\!\!\ln(b+|x|)|\widetilde{u}|^{p}\,{\rm d}x\right)^{\frac{1}{p}}\!\|\widetilde{u}_{n}-\widetilde{u}\|_{\ast,p}\!\left(\int_{\mathbb{R}^{2}}\!\ln(b+|x|)(|\widetilde{u}_{n}|^{p}+|\widetilde{u}|^{p})\,{\rm d}x\right)^{\frac{p-2}{p}}$
$\lesssim\!\left(\int_{B_{R}^{c}}\!\!\ln(b+|x|)|\widetilde{u}|^{p}\,{\rm d}x\right)^{\frac{1}{p}}\left(\|\widetilde{u}_{n}\|_{\ast,p}+\|\widetilde{u}\|_{\ast,p}\right)^{p-1}$
$\leq C\left(\int_{B_{R}^{c}}\!\!\ln(b+|x|)|\widetilde{u}|^{p}\,{\rm d}x\right)^{\frac{1}{p}}\to 0$
as $R\to+\infty$ by Lebesgue’s dominated convergence theorem, since $\|\widetilde{u}\|_{*,p}<+\infty$. Next, again by Hölder’s inequality,
(5.26) $T_{2}$ $\leq\left(\int_{\mathbb{R}^{2}}\ln(b+|x|)|w_{n}|^{r\sigma^{\prime}}{\rm e}^{\alpha\sigma^{\prime}w_{n}^{2}}\,{\rm d}x\right)^{\frac{1}{\sigma^{\prime}}}\!\left(\int_{B_{R}^{c}}\!\!\ln(b+|x|)|\widetilde{u}|^{s\sigma}\,{\rm d}x\right)^{\frac{1}{s\sigma}}$
×$\quad\times\left(\int_{\mathbb{R}^{2}}\ln(b+|x|)|\widetilde{u}_{n}-\widetilde{u}|^{s^{\prime}\sigma}\,{\rm d}x\right)^{\frac{1}{s^{\prime}\sigma}}.$
We first note that
(5.27) $\left(\int_{\mathbb{R}^{2}}\ln(b+|x|)|\widetilde{u}_{n}-\widetilde{u}|^{s^{\prime}\sigma}\,{\rm d}x\right)^{\frac{1}{s^{\prime}\sigma}}\lesssim\|\widetilde{u}_{n}\|+\|\widetilde{u}\|\leq C$
by Proposition 2.6. Moreover, as in (5.25),
(5.28) $\int_{B_{R}^{c}}\!\!\ln(b+|x|)|\widetilde{u}|^{s\sigma}\,{\rm d}x\to 0$
as $R\to+\infty$, again by Proposition 2.6. Let us now focus on the first term in (5.26), which we denote by $T_{2}^{1}$, aiming at an estimate independent of $n$. An $\varepsilon-$Young’s inequality in the exponent yields
(5.29) $\big{(}T_{2}^{1}\big{)}^{\sigma^{\prime}}$ $\lesssim\int_{\mathbb{R}^{2}}\ln(b+|x|)\left(|\widetilde{u}_{n}|^{r\sigma^{\prime}}+|\widetilde{u}|^{r\sigma^{\prime}}\right){\rm e}^{\alpha\sigma^{\prime}\left[(1+\varepsilon)\widetilde{u}_{n}^{2}+C_{\varepsilon}\widetilde{u}^{2}\right]}\,{\rm d}x$
$\leq\left(\int_{\mathbb{R}^{2}}\!\ln(b+|x|)|\widetilde{u}_{n}|^{r\sigma^{\prime}}{\rm e}^{\alpha\sigma^{\prime}\nu(1+\varepsilon)\widetilde{u}_{n}^{2}}\,{\rm d}x\right)^{\frac{1}{\nu}}\!\!\left(\int_{\mathbb{R}^{2}}\!\ln(b+|x|)|\widetilde{u}_{n}|^{r\sigma^{\prime}}{\rm e}^{\alpha\sigma^{\prime}\nu^{\prime}C_{\varepsilon}\widetilde{u}^{2}}\,{\rm d}x\right)^{\frac{1}{\nu^{\prime}}}$
$\quad+\left(\int_{\mathbb{R}^{2}}\!\ln(b+|x|)|\widetilde{u}|^{r\sigma^{\prime}}{\rm e}^{\alpha\sigma^{\prime}\overline{\nu}(1+\varepsilon)\widetilde{u}_{n}^{2}}\,{\rm d}x\right)^{\frac{1}{\overline{\nu}}}\!\!\left(\int_{\mathbb{R}^{2}}\!\ln(b+|x|)|\widetilde{u}|^{r\sigma^{\prime}}{\rm e}^{\alpha\sigma^{\prime}\overline{\nu}^{\prime}C_{\varepsilon}\widetilde{u}^{2}}\,{\rm d}x\right)^{\frac{1}{\overline{\nu}^{\prime}}}$
$=:S_{1}S_{2}+S_{3}S_{4}\,.$
By Theorem 2.3, we first have
(5.30) $S_{4}<+\infty\,.$
Moreover, defining again $v_{n}:=G(\widetilde{u}_{n})$, by Lemma 4.6 and (4.35) we estimate $S_{1}$ as follows:
$S_{1}^{\nu}$ $\leq\frac{{\rm e}^{\alpha\sigma^{\prime}\nu(1+\varepsilon)C_{\varepsilon}t_{\varepsilon}^{2}}}{\tau^{r\sigma^{\prime}/2}}\int_{\mathbb{R}^{2}}\ln(b+|x|)|v_{n}|^{r\sigma^{\prime}}{\rm e}^{\alpha\sigma^{\prime}\nu\frac{(1+\varepsilon)^{2}}{(1-\varepsilon)^{2}}v_{n}^{2}}\,{\rm d}x$
$=C_{\varepsilon}\|\nabla v_{n}\|_{2}^{r\sigma^{\prime}}\!\int_{\mathbb{R}^{2}}\ln(b+|x|)\left(\frac{|v_{n}|}{\|\nabla v_{n}\|_{2}}\right)^{r\sigma^{\prime}}\!{\rm e}^{\alpha\sigma^{\prime}\nu\left(\frac{1+\varepsilon}{1-\varepsilon}\right)^{2}\|\nabla v_{n}\|_{2}^{2}\left(\frac{v_{n}}{\|\nabla v_{n}\|_{2}}\right)^{2}}\,{\rm d}x\,.$
Choose $\alpha>\alpha_{0}$ close to $\alpha_{0}$, $\sigma^{\prime},\,\nu>1$ both close to $1$, and $\varepsilon$ small, so that
$\alpha\sigma^{\prime}\nu\left(\frac{1+\varepsilon}{1-\varepsilon}\right)^{2}\|\nabla v_{n}\|_{2}^{2}<4\pi$
by (4.33), and then $r=p/\sigma^{\prime}$. Since $\|v_{n}\|_{\ast,p}\leq\|\widetilde{u}_{n}\|_{\ast,p}\leq C$, we can apply the weighted Cao’s inequality of Theorem 2.8 and get
(5.31) $S_{1}^{\nu}\leq C(\|v_{n}\|_{\ast,p})\,\|\nabla v_{n}\|_{2}^{p}\leq C$
again by (4.33). It remains only to bound the terms $S_{2}$ and $S_{3}$, which mix $\widetilde{u}_{n}$ and $\widetilde{u}$. On the one hand,
(5.32) $S_{3}^{\overline{\nu}}$$\leq\int_{\{\widetilde{u}\leq\widetilde{u}_{n}\}}\ln(b+|x|)|\widetilde{u}_{n}|^{p}{\rm e}^{\alpha\sigma^{\prime}\overline{\nu}(1+\varepsilon)\widetilde{u}_{n}^{2}}\,{\rm d}x+\int_{\{\widetilde{u}>\widetilde{u}_{n}\}}\ln(b+|x|)|\widetilde{u}|^{p}{\rm e}^{\alpha\sigma^{\prime}\overline{\nu}(1+\varepsilon)\widetilde{u}^{2}}\,{\rm d}x\leq C$
by choosing $\overline{\nu}>1$ close to $1$ and reasoning as for (5.31) for the first term, and by Theorem 2.3 for the second. On the other hand,
(5.33) $S_{2}^{\nu^{\prime}}$ $=\int_{\mathbb{R}^{2}}\ln(b+|x|)|\widetilde{u}_{n}|^{p}{\rm e}^{\alpha\sigma^{\prime}\nu^{\prime}C_{\varepsilon}\widetilde{u}^{2}}\,{\rm d}x$
$\leq C_{\varepsilon}\int_{\{\widetilde{u}\leq 1\}}\ln(b+|x|)|\widetilde{u}_{n}|^{p}\,{\rm d}x+\int_{\{\widetilde{u}>1\}}\ln(b+|x|)|\widetilde{u}_{n}|^{p}\widetilde{u}\,{\rm e}^{\alpha\sigma^{\prime}\nu^{\prime}C_{\varepsilon}\widetilde{u}^{2}}\,{\rm d}x$
$\leq C+\left(\int_{\mathbb{R}^{2}}\ln(b+|x|)|\widetilde{u}_{n}|^{pp^{\prime}}\,{\rm d}x\right)^{\frac{1}{p^{\prime}}}\!\left(\int_{\mathbb{R}^{2}}\ln(b+|x|)\,\widetilde{u}^{p}\,{\rm e}^{\alpha\sigma^{\prime}\nu^{\prime}pC_{\varepsilon}\widetilde{u}^{2}}\,{\rm d}x\right)^{\frac{1}{p}}\leq C\,,$
by Proposition 2.6 and Theorem 2.3. Hence, combining (5.30)-(5.33), we obtain that the term $T_{2}^{1}$ in (5.29) is uniformly bounded. This, together with (5.27) and (5.28), imply that $T_{2}\to 0$ as $R\to+\infty$. Recalling that $T_{1}$ had the same behaviour by (5.25), by (5.24) and (5) this yields
$\int_{\mathbb{R}^{2}}\ln(b+|x|)\left|\widetilde{f}(\widetilde{u}_{n})-\widetilde{f}(\widetilde{u}_{n}-\widetilde{u})\right||\widetilde{u}_{n}-\widetilde{u}|\,{\rm d}x\to 0$
as $n\to+\infty$, which leads to (5.17). By (5.14), this eventually implies both $\|\nabla\widetilde{u}_{n}-\nabla\widetilde{u}\|_{2}\to 0$ and (5.16). By Lemma 3.1(b) then $\|\widetilde{u}_{n}-\widetilde{u}\|_{\ast,p}\to 0$ as $n\to+\infty$ and therefore $\widetilde{u}_{n}\to\widetilde{u}$ in $X$. Since $\mathcal{I}$ is a $C^{1}$-functional, then $\widetilde{u}$ is a weak solution of (1.1), which is nontrivial thanks to (5.8). Since $\widetilde{u}_{n}\geq 0$, by Remark 1.6 the solution $\widetilde{u}$ is positive in $\mathbb{R}^{2}$.
@@
$\int_{\mathbb{R}^{2}}\frac{|\Phi_{u}(x)|}{1+|x|^{2+2s}}\,{\rm d}x$ $\leq\int_{\mathbb{R}^{2}}F(u(y))\left(\int_{\mathbb{R}^{2}}\left|\ln\frac{1}{|x-y|}\right|\frac{1}{1+|x|^{2+2s}}\,{\rm d}x\right)\,{\rm d}y$
$\leq\int_{\mathbb{R}^{2}}F(u(y))\left(\int_{\{|x-y|>1\}}\frac{\ln|x-y|}{1+|x|^{2+2s}}\,{\rm d}x+\int_{\{|x-y|\leq 1\}}\ln\frac{1}{|x-y|}\,{\rm d}x\right)\,{\rm d}y$
$\leq\|F(u(y))\|_{1}\int_{\mathbb{R}^{2}}\frac{\ln(1+|x|)}{1+|x|^{2+2s}}\,{\rm d}x+\|F(u(y))\|_{\ast,1}\int_{\mathbb{R}^{2}}\frac{\,{\rm d}x}{1+|x|^{2+2s}}$
$\quad+\|\ln(\cdot)\|_{L^{1}(B_{1})}\int_{\mathbb{R}^{2}}F(u(y))\,{\rm d}y<+\infty$
@@
$\widetilde{w}_{u}(x)-\Phi_{u}(x)$ $=\int_{\mathbb{R}^{2}}\left(\ln\left(\frac{1+|y|}{|x-y|}\right)-\ln\left(\frac{1}{|x-y|}\right)\right)F(u(y))\,{\rm d}y$
$=\int_{\mathbb{R}^{2}}\ln(1+|y|)F(u(y))\,{\rm d}y<\|F(u)\|_{\ast,1}<+\infty\,,$
@@ Acknowledgements:
The authors are members of Gruppo Nazionale per l’Analisi Matematica, la Probabilità e le loro Applicazioni (GNAMPA) of the Istituto Nazionale di Alta Matematica (INdAM). Federico Bernini is partially supported by INdAM-GNAMPA Project 2024 titled Problemi spettrali e di ottimizzazione di forma: aspetti qualitativi e stime quantitative. Giulio Romani and Cristina Tarsi are partially supported by INdAM-GNAMPA Project 2023 titled Interplay between parabolic and elliptic PDEs and INdAM - GNAMPA Project 2024 titled New perspectives on Choquard equation through PDEs with local sources.